# On the distance to the black hole X-ray binary Swift J1727.8--1613

### v1

@@ 2.1 MeerKAT radio data
All observations were performed with the L-band receiver; two using MeerKAT’s standard “32k” mode, and two using the “32k zoom” mode, hereafter referred to as “32k-S” and “32k-Z” respectively. While each mode contains 32,768 channels, the 32k-Z mode has channel bandwidths that are eight times smaller and thus provides an eight-fold increase in frequency resolution. We alternated our observation scans between J1727 and the phase calibrator, PKS J1733--1304 (J1733 hereafter), with a single scan of the bandpass and flux calibrator J1939--6342 in each observation.
Having obtained MeerKAT observations of J1727 during which it was sufficiently bright at 1.4 GHz (i.e., $\gtrsim\!50$ mJy), we compute H i absorption spectra by processing the radio data to create a radio spectrum that includes the frequency of the H i spectral line, $f_{\textsc{Hi}}=1420.30575177$. We convert frequency to line-of-sight LSR velocity by combining the Doppler shift with the Milky Way rotation curve. We then use the maximum positive or negative velocity observed in the resulting spectra to generate estimates of the kinematic distance via the source code for the Kinematic Distance Calculation Tool [^2] http://www.treywenger.com/kd/, [^3] http://github.com/tvwenger/kd (KDCT; Wenger, 2018).
Figure 1: Our H i absorption spectra along of the line of sight towards J1727. The top two plots are the 32k-S spectra, and the bottom two plots are the 32k-Z spectra, with LSR velocity bin widths that are 5.6 and 0.7 km s-1 wide respectively. The dashed lines indicate the origins of the x- and y-axis. The y-axis represents the residuals after subtracting the continuum emission from the data, with $3\sigma$ uncertainties indicated by the shaded areas, both of which are presented as a percentage of the continuum flux density of the source. The vertical dotted lines indicate the maximum velocity taken as the mid-point of the bin at which significant ($>\!3\sigma$) absorption is observed. In chronological order, these maximum velocities are $19.2\pm 2.8$, $6.1\pm 0.4$, $24.8\pm 2.8$, and $18.0\pm 0.4$ km s-1. The significantly greater SNR of the 14-10-2023 spectrum meant we were able to observe significant ($>\!3\sigma$) H i absorption to greater velocities than observed in the 32k-Z spectra.
@@ 2.1.1 Reference source selection
Due to the lack of bright ($\gtrsim\!50$ mJy) extragalactic background sources in the field i.e., within $1^{\circ}$ of J1727, we used the sufficiently bright ($S_{0}\approx 6$ Jy) phase calibrator J1733 to derive our reference H i absorption spectra. With J1733 located at $(l,b)\approx(12.03^{\circ},10.81^{\circ})$, the two fields are only $3.4^{\circ}$ apart, primarily in Galactic longitude. As J1733 is extragalactic, observations allow us to probe the full set of H i clouds along a nearby line of sight (see also Section 4.1.2 regarding H i scale height).
@@ 2.1.2 Data reduction
We undertook all data reduction on the Ilifu research cloud infrastructure managed by the Inter-University Institute for Data Intensive Astronomy (IDIA) [^4] http://idia.ac.za/ilifu-research-cloud-infrastructure/. To streamline the processing of our H i data, we used the ThunderKAT H i Pipeline [^5] http://github.com/tremou/thunderkat_hi_pipeline.git. Simultaneously, we used CARTA (Cube Analysis and Rendering Tool for Astronomy; Comrie et al., 2021) to interrogate the data.
There are three stages of the pipeline, each with its own bash script that employs several Python scripts.
At a high level, the first stage of the pipeline uses CASA (Common Astronomy Software Applications; CASA Team et al., 2022) to retrieve the data for specified fields from the full observation measurement set and create separate files for each source. The pipeline then undoes previously applied flags to to ensure H i spectral lines are not erroneously flagged as radio-frequency interference. The measurement set for each field is then converted into the fits format required for the Miriad software (Sault et al., 1995) used in the next stage.
@@ 2.2 Hubble Space Telescope near-UV spectroscopy
[…]
These observations provide an alternative and generally more accurate method to estimate $E(B\!-\!V)$ and subsequently the distance to J1727 as detailed in Sections 3.2 and 4.3 respectively.
Figure 2: A consolidated view of the H i spectra for both the target, J1727 (red, left), and our reference source, J1733 (blue, right). We observed no significant ($>\!3\sigma$) absorption outside the velocities chosen as the x-axis range. The y-axis represents the H i absorption percentage. The dashed lines indicate the x- and y-axes origins. The vertical dotted lines indicate the maximum velocity at which significant absorption is observed. For J1727, these are $24.8\pm 2.8$ and $18.7\pm 0.4$ km s-1. For J1733, these are $30.4\pm 2.8$ and $32.7\pm 0.4$ km s-1. The top-left plot is the 32k-S spectrum for J1727 from 14-10-2023, while the bottom-left plot is the mean weighted J1727 spectrum achieved by combining both 32k-Z spectra from Figure 1. We generated the 32k-Z (32k-S) spectra for J1733 from observations on 06-09-2023 (14-10-2023) with the peak J1733 flux density measured to be $6.1\pm 0.1$ Jy ($6.0\pm 0.1$ Jy), providing the significantly greater SNR compared to that of the J1727 observations.
Figure 3: The near-UV spectrum of J1727 as seen by HST/STIS (black), with a reddened power-law fit to the data (green) (Castro Segura et al. in prep.) The uncertainty around the fit is indicated as a shaded region. The vertical shaded regions were masked during the fit to avoid Fe and Mg lines that may influence the fit.
@@ 3.1.1 High resolution 32k zoom mode
Both H i spectra from our two 32k-Z observations are displayed in the bottom two plots of Figure 1. Significant ($>\!3\sigma$) H i absorption towards J1727 is observed out to an estimated maximum LSR velocity $v_{\rm{LSR}}=18.7\pm 0.4$ km s-1 as shown in the inset for the the mean weighted 32k-Z spectrum at the bottom-left of Figure 2, using half the bin width as the velocity uncertainty. The maximum absorption is observed to be $\approx 90\%$.
@@ 3.2 Near-UV
We determined the line of sight extinction to J1727 using the Cardelli et al. (1989) reddening law to fit a reddened power-law representing the outer accretion disk to the near-UV spectrum as displayed in Figure 3. To estimate the errors we performed a Monte Carlo simulation yielding $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$, corresponding to $A_{V}=R_{V}\,E(B\!-\!V)\simeq 1.00$ for $R_{V}=3.1$.
@@ 4.1 Kinematic distance constraint
We used our maximum H i absorption velocity and its uncertainty as inputs to the KDCT (Wenger et al., 2018), using their Method C and applying the revised solar motion parameters from Reid et al. (2014) and the rotation curve from Reid et al. (2019). The total revised LSR velocity uncertainty includes measurement and systematic uncertainties such as non-circular streaming motions.
@@
The quantities of $d_{\rm{tan}}$ and $v_{\rm{LSR,tan}}$ depend only on the source location. Therefore, we used $(l,b)=(8.641502^{\circ},10.254899^{\circ})$ for J1727 to estimate $d_{\rm{tan}}=8.20\pm 0.03$ kpc and $v_{\rm{LSR,tan}}=93.9\pm 3.3$ km s-1. As $d_{\rm{near}}$ depends on the input velocity, we use our higher-SNR result of $24.8\pm 2.8$ km s-1 from our 32k-S spectra to estimate $d_{\rm{near}}=3.6\pm 0.3$ kpc.
@@ 4.1.1 Caveats
Regions within 15 degrees in Galactic longitude from the GC feature increased H i emission from the GC (Kalberla & Kerp, 2009), which leads to higher sky temperatures around the H i line and hence increased uncertainty in each spectral channel. At these longitudes, the motion of objects intersected by the line of sight is mostly perpendicular to the line of sight. Distances are thus inferred from a smaller spread in circular rotation velocities and subject to larger uncertainties, and so this region was excised from the study of Wenger et al. (2018).
More recently, Hunter et al. (2024) conducted numerical 2D hydrodynamical simulations to account for potential causes of Milky Way deviations from axisymmetry and gas cloud deviations from the circular rotation curve. The authors categorise these deviations into: (i) random fluctuations around the average streaming motions that do not change the average velocity; and (ii) systematic changes in streaming velocity due to non-axisymmetry, such as spiral arms and the Galactic bar. The authors then define regions of their simulated Milky Way Galaxy where the discrepancy between the kinematic and true distance is significant ($>\!27\%$). Within $52\!<\!d\!<\!$ 5 kpc, the longitude for J1727 appears to correspond to a median absolute relative kinematic distance error of $|d_{\rm{k}}-d_{\rm{true}}|/d_{\rm{true}}\approx 63\%$. For $105\!<\!d\!<\!$ 10 kpc, this error reduces to approximately $12\pm 8\%$. We use the greater 63% error above to expand the $1\sigma$ uncertainty on our measured kinematic distance, which becomes $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc. Having used 2D simulations, Hunter et al. (2024) assume that the gas is integrated along the $z$-axis (vertically), and that the acceleration of the gas due to the Galactic potential is computed as if the gas lies in the GP with Galactic elevation, $z$, equal to zero. Therefore, with a Galactic latitude of $b\approx 10.25^{\circ}$, the total error due to the Galactic longitude of J1727 may not amount to the full 63%. Nevertheless, we observe that minor changes in Galactic longitudes close to the GC have major impacts on the values of $d_{\rm{near}}$ calculated using the KDCT. Specifically, J1733 has a maximum H i absorption velocity that is more than 30% greater than that of J1727, however its larger Galactic longitude of $l\approx 12^{\circ}$ results in a similar predicted value of $d_{\rm{near}}$.
Secondly, the kinematic distance method works best for sources located in the GP where $b\!\approx\!0$. Wenger et al. (2018) assumed a latitude of $b\!=\!0$ and only use latitude to correct the LSR velocity with updated solar motion parameters. High Galactic latitudes correspond to greater Galactic elevations where less gas and other matter reside. Observed absorption features are primarily due to the gas clouds that are closer to us, as evidenced by the absence of detectable H i absorption out to the tangent point towards J1733. The impact of the Galactic latitude for J1727 will therefore lead to, if anything, an underestimation of the distance.
More conservatively estimating the lower limit of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc by accounting for longitudinal effects may therefore be reasonable, given the high Galactic latitude of J1727 and thus diminishing H i density along the lines of sight towards J1727 and J1733.
@@ 4.1.2 Scale height of Galactic H I
H i absorption is not observed to the tangent point velocity of $93.9\pm 3.3$ km s-1 in any of our spectra, only up to a maximum velocity $32.7\pm 0.4$ km s-1 in the direction of J1733. This suggests that the line of sight has not intersected H i clouds at greater distances due to $z$ increasing and H i density decreasing with distance. The distance lower bound of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc corresponds to a GP elevation of $z\approx 650\pm 410$ pc.
@@ 4.1.3 Feasibility of H I absorption XRB distances
4.2 Comparison with independent distance estimates
Optical techniques like those introduced in Sections 1.1.2 and 1.1.3 have long been used to estimate distances to XRBs by using the distance modulus. MS24a used donor star magnitudes in conjunction with various relations in the literature to derive values for the parameters in Equation 1, and calculated the weighted mean of the resulting distances to be $d=2.7\pm 0.3$ kpc. These distance estimates, and therefore the consolidated weighted mean value, were improved upon by MS24b using direct measurement of the orbital period, $P_{\rm{orb}}$. MS24b also report the best fitting spectral type template of K3/4V for a donor star that is partially veiled by the accretion disk. This veiling factor of $X_{\rm{veil}}=74\%$ is defined as the ratio of fluxes of non-stellar origin to the total emitted light (in the $r$-band). The authors combined this with their measured mean $r$-band magnitude of $r=19.55\pm 0.14$ to derive the donor star apparent $r$-band magnitude $m_{r}=m_{\rm{star}}=21.13\pm 0.05$. The K4V template also allows the authors to revise the absolute $r$-band magnitude to $M_{r}=6.6\pm 0.5$. These efforts resulted in an updated distance of $d=3.4\pm 0.3$ kpc.
Most existing distance estimates (MS24a,b) relied on inference of $E(B\!-\!V)$. One can then derive the $r$-band extinction using $A_{r}=2.271\,E(B\!-\!V)$ (for Pan-STARRS1; Schlafly & Finkbeiner, 2011) for insertion into Equation 1.
The distance estimation methods employed by MS24a and MS24b include the relations between:
  1. (1)
    The interstellar Ca ii doublet (H and K) and the distance to early-type stars per Megier et al. (2009). As reported by MS24a and MS24b, this yields $d_{1}=3.2\pm 0.6$ kpc. The relation only applies to objects within a few hundred pc from the GP, which is unlikely for J1727 given our discussion of GP elevation in Section 4.1.2.
  2. (2)
    The equivalent width of the interstellar line K i 7699 Å and $E(B\!-\!V)$ per Munari & Zwitter (1997). MS24a’s Monte Carlo analysis ($10^{5}$ samples) gave $E(B\!-\!V)=0.9\pm 0.3$, which MS24b reused to estimate $d_{2}=3.2\pm 1.3$ kpc.
  3. (3)
    The equivalent width of the diffuse interstellar band at 8621 Å  and $E(B\!-\!V)$ per Wallerstein et al. (2007). Using a similar method to (2), MS24a calculated $E(B\!-\!V)=0.8\pm 0.3$, leading MS24b to calculate $d_{3}=3.5\pm 1.5$ kpc.
  4. (4)
    The hydrogen column density $N_{\rm{H}}$ and $V$-band extinction $A_{V}$ per Güver & Özel (2009): $N_{\rm{H}}(\rm{cm}^{-2})=(2.21\pm 0.09)\times 10^{21}\mathit{A_{V}}$. X-ray studies provided $N_{\rm{H}}=(3.2\pm 0.9)\times 10^{21}\rm{cm}^{-2}$ (O’Connor et al., 2023; Draghis et al., 2023). With $A_{V}=3.1\,E(B\!-\!V)$ (Savage & Mathis, 1979; Fitzpatrick, 2004), MS24a estimated $E(B\!-\!V)=0.47\pm 0.13$, which for MS24b yielded $d_{4}=4.8\pm 1.4$ kpc.
To investigate the impact of the significant systematic uncertainties inherent in these relations on the final distance estimates, we employed Monte Carlo techniques to resample each of the parameters involved in these calculations within their uncertainties $10^{5}$ times, assuming normal distributions for all quantities. We find positively skewed probability density functions for the distances resulting from these methods and that MS24b likely give mean distance values.
4.3 Updated $E(B\!-\!V)$ constraints and resulting distance

### v2

@@ 2.1 MeerKAT radio data
All observations were performed with the L-band receiver; two using MeerKAT’s standard “32k” mode, and two using the “32k zoom” mode, hereafter referred to as “32k-S” and “32k-Z” respectively. While each mode contains 32,768 channels, the 32k-Z mode has channel bandwidths that are eight times smaller and thus provides an eight-fold increase in frequency resolution. We alternated our observation scans between J1727 and the phase calibrator, PKS J1733-1304 (J1733 hereafter), with a single scan of the bandpass and flux calibrator J1939-6342 in each observation.
Having obtained MeerKAT observations of J1727 during which it was sufficiently bright (i.e., $\gtrsim\!50$ mJy) at 1.4 GHz, we compute H i absorption spectra by processing the radio data to create a radio spectrum that includes the frequency of the H i spectral line, $f_{\textsc{Hi}}=1420.30575177$ MHz. We convert Doppler shifts in frequency to line-of-sight LSR velocities. We then compare the maximum positive or negative velocity observed in the resulting spectra with the Milky Way rotation curve to generate estimates of the kinematic distance via the source code for the Kinematic Distance Calculation Tool [^2] http://www.treywenger.com/kd/, [^3] http://github.com/tvwenger/kd (KDCT; Wenger, 2018).
@@ 2.1.1 Reference source selection
Due to the lack of bright ($\gtrsim\!50$ mJy) extragalactic background sources in the field i.e., within $1^{\circ}$ of J1727, we used the bright ($S_{0}\approx 6$ Jy) phase calibrator J1733 to derive our reference H i absorption spectra. With J1733 located at $(l,b)\approx(12.03^{\circ},10.81^{\circ})$, the two fields are only $3.4^{\circ}$ apart, primarily in Galactic longitude. As J1733 is extragalactic, observations allow us to probe the full set of H i clouds along a nearby line of sight (see also Section 4.1.2 regarding H i scale height).
Figure 1: Our H i absorption spectra along of the line of sight towards J1727. The top two plots are the 32k-S spectra, and the bottom two plots are the 32k-Z spectra, with LSR velocity bin widths that are 5.6 and 0.7 km s-1 wide respectively. The dashed lines indicate the origins of the $x$- and $y$-axes. The $y$-axis represents the residuals after subtracting the continuum emission from the data, with $3\sigma$ uncertainties indicated by the shaded areas, both of which are presented as a percentage of the continuum flux density of the source. The vertical dotted lines indicate the maximum velocity taken as the mid-point of the bin at which significant ($3\sigma$) absorption is observed. In chronological order, these maximum velocities are $19.2\pm 2.8$, $6.1\pm 0.4$, $24.8\pm 2.8$, and $18.0\pm 0.4$ km s-1. The significantly greater SNR of the 14-10-2023 spectrum meant we were able to observe significant ($3\sigma$) H i absorption to greater velocities than observed in the 32k-Z spectra.
@@ 2.1.2 Data reduction
We undertook all data reduction on the Ilifu research cloud infrastructure managed by the Inter-University Institute for Data Intensive Astronomy (IDIA) [^4] http://idia.ac.za/ilifu-research-cloud-infrastructure/. To streamline the processing of our H i data, we used the ThunderKAT H i Pipeline [^5] http://github.com/tremou/thunderkat_hi_pipeline.git. Simultaneously, we used CARTA (Cube Analysis and Rendering Tool for Astronomy; Comrie et al., 2024) to interrogate the data.
The ThunderKAT H i Pipeline has three stages, each with its own bash script that employs several Python scripts.
At a high level, the first stage of the pipeline uses CASA (Common Astronomy Software Applications; CASA Team et al., 2022) to retrieve the data for specified fields from the full observation measurement set and create separate files for each source. The pipeline then undoes previously applied flags to ensure H i spectral lines are not erroneously flagged as radio-frequency interference. The measurement set for each field is then converted into the fits format required for the Miriad software (Sault et al., 1995) used in the next stage.
@@ 2.2 Hubble Space Telescope near-UV spectroscopy
We obtained high resolution near-UV spectroscopy with the Space Telescope Imaging Spectrograph (STIS; Woodgate et al., 1998) aboard the Hubble Space Telescope (HST) in early October 2023 ($\rm MJD\sim 60219$) during the outburst (program ID 16489; Castro Segura et al., 2020). We used E230M gratings with 200 s followed by 220 s exposures at the central wavelengths 1978 Å and 2707 Å respectively to cover the region $\lambda\lambda\simeq 1800\text{--}3200$ Å with a resolving power of R=30, [^0] R=30,\!000. The data were reduced using the HST pipeline calstis [^6] Provided by The Space Telescope Science Institute
(https://github.com/spacetelescope)
@@ 3.1 Radio
Figure 2: A consolidated view of the H i spectra for both the target, J1727 (red, left), and our reference source, J1733 (blue, right). We observed no significant ($3\sigma$) absorption outside the velocities chosen as the $x$-axis range. The $y$-axis represents the H i absorption percentage. The dashed lines indicate the $x$- and $y$-axes origins. The vertical dotted lines indicate the maximum velocity at which significant absorption is observed. For J1727, these are $24.8\pm 2.8$ and $18.7\pm 0.4$ km s-1. For J1733, these are $30.4\pm 2.8$ and $32.7\pm 0.4$ km s-1. The top-left plot is the 32k-S spectrum for J1727 from 14-10-2023, while the bottom-left plot is the mean weighted J1727 spectrum achieved by combining both 32k-Z spectra from Figure 1. We generated the 32k-Z (32k-S) spectra for J1733 from observations on 06-09-2023 (14-10-2023) with the peak J1733 flux density measured to be $6.1\pm 0.1$ Jy ($6.0\pm 0.1$ Jy), providing the significantly greater SNR compared to that of the J1727 observations.
@@ 3.1.1 High resolution 32k zoom mode
Both H i spectra from our two 32k-Z observations are displayed in the bottom two plots of Figure 1. Significant ($3\sigma$) H i absorption towards J1727 is observed out to an estimated maximum LSR velocity $v_{\rm{LSR}}=18.7\pm 0.4$ km s-1 as shown in the inset for the the mean weighted 32k-Z spectrum at the bottom-left of Figure 2, using half the bin width as the velocity uncertainty. The maximum absorption is observed to be $\approx 90\%$.
@@ 3.1.3 Reference source
Figure 3: The near-UV spectrum of J1727 as seen by HST/STIS (black), with a reddened power-law fit (green) to the data (Castro Segura et al. in prep.) The uncertainty around the fit is indicated as a shaded region. The vertical shaded regions were masked during the fit to avoid Fe and Mg lines that may influence the fit.
@@ 3.2 Near-UV
We determined the line of sight extinction to J1727 using the Cardelli et al. (1989) reddening law to fit a reddened power-law representing the outer accretion disk to the near-UV spectrum as displayed in Figure 3. To estimate the errors we performed a Monte Carlo simulation yielding $E(B\!-\!V)=0.37\pm 0.01\,({stat})\pm 0.025\,({sys})$, corresponding to $A_{V}\simeq 1.15$ per Equation 2 with $R_{V}=3.1$. Given we use $r$-band magnitudes when calculating Equation 1, we derive the $r$-band extinction using $A_{r}=2.271\,E(B\!-\!V)=0.84\pm 0.06$ (for Pan-STARRS1; Schlafly & Finkbeiner, 2011).
@@ 4.1 Kinematic distance constraint
We used our maximum H i absorption velocity and its uncertainty as inputs to the KDCT (Wenger et al., 2018), using their Method C and applying the revised solar motion parameters from Reid et al. (2014b) and the rotation curve from Reid et al. (2019). The total revised LSR velocity uncertainty includes measurement and systematic uncertainties such as non-circular streaming motions.
@@
The quantities $d_{\rm{tan}}$ and $v_{\rm{LSR,tan}}$ depend only on the source location. For J1727, we therefore estimate $d_{\rm{tan}}=8.20\pm 0.03$ kpc and $v_{\rm{LSR,tan}}=93.9\pm 3.3$ km s-1. As $d_{\rm{near}}$ depends on the input velocity, we use our higher-SNR result of $24.8\pm 2.8$ km s-1 from our 32k-S spectra to estimate $d_{\rm{near}}=3.6\pm 0.3$ kpc.
@@ 4.1.1 Caveats
Regions within $15^{\circ}$ in Galactic longitude from the GC feature increased H i emission from the GC (Kalberla & Kerp, 2009), which leads to higher sky temperatures around the H i line and hence increased uncertainty in each spectral channel. At these longitudes, the motion of objects intersected by the line of sight is mostly perpendicular to the line of sight. Distances are thus inferred from a smaller spread in circular rotation velocities and subject to larger uncertainties, and so this region was excised from the study of Wenger et al. (2018).
More recently, Hunter et al. (2024) conducted numerical 2D hydrodynamical simulations to account for potential causes of Milky Way deviations from axisymmetry and gas cloud deviations from the circular rotation curve. The authors categorise these deviations into: (i) random fluctuations around the average streaming motions that do not change the average velocity; and (ii) systematic changes in streaming velocity due to non-axisymmetry, such as spiral arms and the Galactic bar. The authors then define regions of their simulated Milky Way Galaxy where the discrepancy between the kinematic and true distance is significant ($27\%$). Within 2<d<52\!<\!d\!<\!5 kpc, the longitude for J1727 appears to correspond to a median absolute relative kinematic distance error of $|d_{\rm{k}}-d_{\rm{true}}|/d_{\rm{true}}\approx 63\%$. For 5<d<105\!<\!d\!<\!10 kpc, this error reduces to approximately $12\pm 8\%$. We use the greater 63% error above to expand the $1\sigma$ uncertainty on our measured kinematic distance, which becomes $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc.
We observe that minor changes in Galactic longitudes close to the GC have major impacts on the values of $d_{\rm{near}}$ calculated using the KDCT. Specifically, J1733 has a maximum H i absorption velocity that is more than 30% greater than that of J1727, however its larger Galactic longitude of $l\approx 12^{\circ}$ results in a similar predicted value of $d_{\rm{near}}$.
Our conservative lower limit of $3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc by accounting for longitudinal effects may therefore be reasonable, given the high Galactic latitude of J1727 and thus diminishing H i density along the lines of sight towards J1727 and J1733.
@@ 4.1.2 Scale height of Galactic H I
The kinematic distance method works best for sources located in the GP where $b\!\approx\!0$. Wenger et al. (2018) assumed a latitude of b=0b\!=\!0 and only use latitude to correct the LSR velocity with updated solar motion parameters. Having used 2D simulations, Hunter et al. (2024) assumed that the gas is integrated along the $z$-axis (vertically), and that the acceleration of the gas due to the Galactic potential is computed as if the gas lies in the GP with Galactic elevation, $z$, equal to zero. High Galactic latitudes correspond to greater Galactic elevations where less gas and other matter reside. Observed absorption features are primarily due to the gas clouds that are closer to us, as evidenced by the absence of detectable H i absorption out to the tangent point towards J1733. The impact of the Galactic latitude for J1727 will therefore lead to, if anything, an underestimation of the distance.
We do not observe H i absorption to the tangent point velocity of $93.9\pm 3.3$ km s-1 in any of our spectra; only up to a maximum velocity $32.7\pm 0.4$ km s-1 in the direction of J1733. This suggests that the line of sight has not intersected H i clouds at greater distances due to $z$ increasing and H i density decreasing with distance. The distance lower bound of $d_{\rm{near}}=3.6\pm 0.3\,({stat})\pm 2.3\,({sys})$ kpc corresponds to a GP elevation of $z\approx 650\pm 410$ pc.
@@ 4.1.3 Feasibility of H I absorption XRB distances
4.2 E(BV)E(B\!-\!V) distance constraint
4.2.1 E(BV)E(B\!-\!V) validation